# Rewarding Curse: Analyze and Mitigate Reward Modeling Issues for LLM Reasoning

### v1

@@ 4.2 Discriminative Experiments on RM
Figure 6: The number of times the model’s responses change from correct to incorrect as $N$ changes.
Statistical Analysis
We further record the number of times the answer changes from correct to incorrect and report the results in Figure 6. All of the methods tend to make more incorrect transitions as $N$ increases. This demonstrates that the model tends to make more wrong differentiation under a higher sampling number. Besides, compared to SC, RM-based methods perform worse (have higher counts in Figure 6), which highlights the introduction of reward models leads to more wrong selection.
4.3 Negative Examples Analysis
Hypothesis on the Issue
Why does the reward model perform worse as the sampling number grows? Reflecting on its training process (Wang et al., 2024a; Liu et al., 2024a; Wang et al., 2024b), the training data primarily consists of paired responses (i.e., a correct one and an incorrect one). The limited number of samples covers only a restricted response distribution. Therefore, we hypothesize that: As $N$ increases, a significant number of low-frequency samples, which deviate from the distribution of the training data, are generated. The reward model struggles to generalize to these unseen samples, resulting in higher scores being assigned to incorrect responses within them.
(a) BoN
(b) MCTS
Figure 7: Frequency statistics of the highest-scored negative samples.
Experimental Setup
To validate our hypothesis, we conduct a statistical analysis of negative examples. For each question, We select the incorrect response with the highest RM score and then calculate the frequency of its answer across all samples. Then, we analyze the distribution of this frequency across 500 questions.
Main Results
The results depicted in Figure 7 illustrate that the RM exhibits an inverse long-tail phenomenon during scoring wrong responses. On the vast majority of questions, incorrect answers with low occurrence frequencies ($frequency$ < 5 in Figure 7) receive the highest scores among all negative examples. Conversely, incorrect answers with high occurrence frequencies rarely achieved the highest scores ($counts$ < 10 in Figure 7). The results effectively substantiate our earlier hypothesis: (Cl.2) Due to the limited distribution covered by the RM’s paired training data, it struggles to correctly generalize to incorrect responses with low occurrence frequencies, making it increasingly difficult to distinguish incorrect responses from correct ones as $N$ grows.
5 Search Diversity: RM performs worse on high-diversity distributions
The final influencing factor we investigate is the search parameters $\Phi$. It encompasses various parameters within the search strategy, which are primarily utilized to control the diversity of the policy model’s search. In BoN inference, we study the influence of the temperature $T$ on RM’s performance (§5.1). Then, in MCTS, we mainly investigate the impact of tree structures on RM (§5.2).
(a) Qwen2.5-3B
(b) Llama3.1-8B
Figure 8: Performance of BoN inference across different sampling temperatures.
(a) Width of ORM
(b) Width of PRM
(c) Max depth of ORM
(d) Max depth of PRM
Figure 10: MCTS inference performance under different tree structures.
Figure 11: Main process of our OCTS algorithm.
5.1 Search Diversity in BoN
Overall Performance
When we directly sample responses from LLMs (like SC and BoN), the temperature $T$ is a key parameter to control the search diversity. Here, we traverse different temperatures to study the performance variations of BoN and demonstrate the results in Figure 8. For both policy models, as the temperature increases, the performance of BoN continues to decline. In contrast, SC and Oracle (i.e. coverage) exhibit more stable performance, showing a declining trend only when the sampling temperature exceeds a relatively high threshold ($T$ > 0.9 in Figure 8). This indicates that RM is more sensitive to sampling diversity compared to the policy model. Higher diversity makes it challenging for the RM to distinguish between positive and negative responses.
Statistical Analysis
We further conduct statistical analyses to uncover the reasons for this issue. For each $T$, we calculate the information entropy of incorrect answers across 16 samplings and report the distribution over 200 questions in Figure 9. As the temperature rises, the entropy for both models shows a gradually increasing trend, hence, the distribution of these negative samples becomes more random. This indicates that the policy model generates a greater number of low-frequency incorrect answers at higher temperatures. According to Cl.2 in §4.3, RM struggles to differentiate these negative examples from correct ones, leading to lower inference accuracy. This result not only elucidates the reasons behind the subpar performance of BoN under high diversity conditions but also further corroborates the inverse long-tail phenomenon of the RM.
5.2 Search Diversity in MCTS
Experimental Setup
In the MCTS algorithm, the main parameter governing search diversity is the tree structure, which is defined by two key parameters: width and maximum depth. The width refers to the number of child nodes at each node. A larger width corresponds to a broader search space in the exploration process. The depth represents the length of the longest path from the root node to the leaf node. A greater depth suggests that the model can explore more intermediate states along a single branch. We set the default width and maximum depth to 5 and compare the reasoning performance of MCTS under different parameters.
Main Results
The results of the comparison are illustrated in Figure 10. From the figure, we can conclude that: (1) For width, the best performance is observed at intermediate values (width = 5), too high widths lead to a decline in performance. This indicates that MCTS should prevent the majority of rollouts from being used solely for exploration. (2) For max depth, the best performance is achieved under settings with a lower value (i.e., max depth = 3 or 5). This suggests that, for MCTS, an excessive number of intermediate states (max depth > 5) can impair performance. In practical searches, the optimal number of intermediate states does not align with the steps a human would take to solve the problem. To sum up, excessive diversity from both width and depth can degrade the RM’s performance. Thus, we have: (Cl.3) Regardless of the search strategy, it is essential to limit the diversity of the sampling distribution to maintain the optimal performance of the RM. We also study the diversity of MCTS based on the explore weight, and the results are consistent (see Appendix E).
6 Application: OCTS boosts RM-based reasoning performance
6.1 Our Method
@@
In the preceding sections, we uncover key patterns that affect the RM’s performance and identify serval issues in RM-based reasoning. To mitigate these issues, we propose a novel RM-based inference algorithm called Optimal Clustering Tree Search (OCTS), as an application of our findings. Figure 11 demonstrates the main process of our algorithm, which comprises three key steps:
Exploration
We initially employ the BoN strategy for exploration, directly generating complete reasoning paths. If the number of answers is below a certain threshold, it indicates that the question is simple (see Appendix C for details). In such case, the algorithm stops directly, selecting the path with the highest score from the majority-voted answers as the final output. This not only reduces the additional inference costs but also avoids the RM’s detrimental impact on the reasoning performance for simple questions. Thus, this early stopping mechanism can mitigate the issue in Cl.1.
Selection
In this step, we cluster all paths based on their answers and sum the reward scores of all paths within the same cluster to derive a cluster score. For example, in Step 2 of Figure 11, we have two clusters, one with a score of 1.3 (0.6 + 0.7 = 1.3) and the other with a score of 0.8. Then, using these cluster scores as the criterion, we select the top-k clusters and choose the best path from each of them. Therefore, we can use answer frequency as another key factor in path scoring to reduce the scores of low-frequency negative examples, mitigating the issue in Cl.2.
Expansion
After selecting k paths, we expand the tree based on them. Specifically, in the n-th round, we use the first n solution steps of each selected path as the intermediate states (see intermediate solution step in Step 3 of Figure 11) to generate new children nodes. All remaining steps are generated directly to reduce the intermediate nodes in the tree. By controlling the complexity of the tree, we can reduce the negative impact of excessive diversity on the RM, as described in Cl.3.
We repeat steps 2 and 3 until the tree reaches the maximum depth and select the best path in the last round as the final output.
Methods GSM8k MATH
ORM PRM ORM PRM
CoT 78.2 78.2 46.2 46.2
Self-Consistency 82.8 82.8 64.2 64.2
Best-of-N 83.0 86.8 64.6 61.2
BoN Weighted 83.4 86.2 66.6 59.8
MCTS 91.8 94.8 66.6 70.6
Beam Search - 94.6 - 72.6
Ours 90.8 95.8 69.8 75.6
@@
Table 1: Performance comparison in main experiments, the best results are highlighted in bold.
@@ 6.2 Main Experiments
We compare the reasoning performance of our method with other advanced baselines, including: CoT (Wei et al., 2022), Self-Consistency (Wang et al., 2023), Best-of-N, BoN Weighted (Snell et al., 2024), MCTS (Hao et al., 2023) and Beam Search (Snell et al., 2024). For datasets, in addition to MATH-500 (Hendrycks et al., 2021; Lightman et al., 2024), we also validate our methods on GSM8k (Cobbe et al., 2021). For models, we select Qwen2.5-3B and Llama3.1-8B as the policy model, while using Skywork-Llama-3.1-8B (ORM) and Skywork-o1-PRM-Qwen-2.5-7B (PRM) as the reward model. We present more details of the implementation in Appendix F.
@@
We demonstrate the result on Qwen-2.5-3B in Table 1 (more results, including more datasets, more models, and the ablation study are presented in Appendix G), from which we can get the following conclusions: (1) Our OCTS method enables more effective utilization of the reward model in reasoning tasks. Compared to other RM-based inference methods, it effectively improves reasoning accuracy across different datasets and models, with an improvement of up to 3.2%. (2) The findings drawn from the former analysis are reasonable. Given that our method is an application derived from the analytical conclusions, its superior performance can further substantiate the correctness of our earlier conclusions.
7 Related Works
Inference-time Scaling Technique
@@ 7 Related Works
The emergence of o1-like models has demonstrated that inference-time scaling effectively enhances LLM’s reasoning abilities (OpenAI, 2024; DeepSeek-AI et al., 2025; Zeng et al., 2024; Zhao et al., 2024). Existing work primarily follows two approaches: optimizing the strategy for LLMs to search for answers (Hao et al., 2023; Snell et al., 2024; Bi et al., 2024; Qi et al., 2024) or improving the reward model’s ability to evaluate response quality (Wang et al., 2024b; Zhang et al., 2024; Setlur et al., 2024). However, most studies explore these two approaches separately, with limited research analyzing the impact of search factors (e.g. sampling numbers, search diversity) on RM performance. Our work addresses this gap and proposes a new search strategy to mitigate RM’s deficiencies.
@@
The reward model plays a crucial role in complex reasoning tasks of LLMs (Zeng et al., 2024; Setlur et al., 2024; Wang et al., 2024b). Existing works mainly investigate the RM from two perspectives: evaluation and optimization. For the former, researchers design various datasets to evaluate the RM’s ability to distinguish between positive and negative responses (Lambert et al., 2024; Liu et al., 2024b; Zheng et al., 2024). For the latter, researchers focus on the training phase, improving the RM’s ability by synthesizing high-quality data (Wang et al., 2024b; Liu et al., 2024a) or optimizing the training algorithm (Zhang et al., 2024; Ankner et al., 2024; Lou et al., 2024). There is a lack of in-depth analysis of the potential issues RM faces during inference, as well as methods to optimize RM’s performance in the inference stage. Our work addresses the gaps left by these related studies.

### v2

@@ 3.4 Sampling Number: RM struggles to distinguish low-frequency negatives
Figure 4: The number of times the model’s selection changes from correct to incorrect.
Figure 5: Frequency statistics of the highest-scored negative responses in BoN.
Performance Gap between Accuracy and Coverage
Recent studies (Brown et al., 2024) show that the coverage of correct answers by LLMs increases with the number of samples, while accuracy plateaus after a small $n$ (see Appendix G for experimental details). Given that recall steadily improves, we suggest that the accuracy bottleneck is likely a result of the RM making more misclassifications as $n$ increases. To investigate this, we first conduct a case study in which we randomly select questions and examine the RM’s selection accuracy at different $n$ (see Appendix H for details). The results indicate that, in some cases, the RM assigns the highest score to incorrect responses generated at higher $n$, replacing originally correct answers with incorrect ones. Based on the observation, we further record the number of instances in which the selected answer transitions from correct to incorrect and present the results in Figure 5. All methods exhibit a tendency for more incorrect transitions as $n$ increases. Compared to SC, RM-based inference methods show higher transition counts in Figure 5, which suggests that incorporating reward models introduces more incorrect selections.
Inverse Long-tail Phenomenon
Why does the reward model perform worse as the sampling number grows? Reflecting on its training process (Wang et al., 2024a; Liu et al., 2024a; Wang et al., 2024b), the training data primarily consists of paired responses (i.e., a correct one and an incorrect one). These pairs represent a constrained subset of the response space. We hypothesize that as $n$ grows, more low-frequency responses (those outside the training distribution) are sampled. The reward model struggles to generalize to these unfamiliar inputs, leading to incorrect responses occasionally receiving higher scores. To validate this hypothesis, we perform a statistical analysis of negative responses. For each question, we select the incorrect response with the highest RM score and compute the frequency of its answer across all samples. As shown in Figures 5 and 23, the RM displays an inverse long-tail phenomenon when scoring incorrect responses. For most questions, the top-scoring incorrect answers tend to have very low frequencies (frequency <5<5 in Figure 5). Conversely, incorrect answers with high occurrence frequencies rarely achieved the highest scores. These findings support our hypothesis: (Cl.2) RMs struggle to correctly score incorrect responses with low occurrence frequencies, making it difficult to distinguish incorrect responses from correct ones as $n$ grows.
3.5 Search Parameters: RM performs worse on high-diversity distributions
Search Diversity in BoN
The final influencing factor we investigate is the search parameters $\Phi$, which are primarily utilized to control the diversity of the policy model’s search. For the BoN method, the temperature $T$ is the key parameter controlling the search diversity. We sweep $T$ and analyze its influence on the performance, as shown in Figure 7 and 24. For both policy models, BoN performance consistently degrades with increasing $T$, while SC and Oracle (i.e., coverage) remain stable except at high temperatures ($T>0.9$ in Figure 7). These results indicate that RM is more sensitive to sampling diversity than the policy model. Higher diversity makes it challenging for the RM to distinguish between positive and negative responses. To better understand this issue, we perform additional statistical analyses in Appendix I, which suggest that higher sampling temperatures cause the policy model to produce more low-frequency incorrect responses, thereby degrading discriminative accuracy.
Figure 6: BoN performance across different temperatures (Qwen2.5-3B).
(a) Tree width
(b) Tree depth
Figure 7: MCTS performance under different tree structures (ORM).
Search Diversity in MCTS
In the MCTS algorithm, search diversity is primarily governed by the tree structure, determined by two key parameters: width and depth. The width refers to the number of child nodes at each node, whereas the depth denotes the length of the longest path from the root to a leaf node. A larger width indicates a broader search space during exploration, while a greater depth implies the model can traverse more intermediate states along a single trajectory. We evaluate MCTS performance under varying settings and present the results in Figure 7 and 27 . The findings reveal: (1) For width, the best performance is observed at intermediate values (width = 5), too high widths lead to a decline in performance. (2) For depth, the best performance is achieved under settings with a lower value (e.g., depth = 3 or 5). These suggest that in MCTS, exploring too many intermediate states can harm performance. Notably, the optimal number of intermediate steps in search does not necessarily align with the number of steps a human would take to solve the same problem. We also analyze the impact of exploration weight on the diversity of MCTS, with consistent findings (see Appendix J). In summary, excessive diversity, such as width, depth, or temperature, can impair the performance of the reward model. Thus, we conclude: (Cl.3) During inference, it is essential to constrain the diversity of the sampling distribution to maintain the optimal performance of the RM.
4 Mitigating RM-based Inference Issues
4.1 Our Methodology
Figure 8: Main process of our CRISP method.
@@ 4.1 Our Methodology
In the preceding sections, we uncover key patterns that affect the RM’s performance and identify serval issues in RM-based reasoning. To mitigate these issues, we propose a novel RM-based inference algorithm called Clustered Reward Integration with Stepwise Prefixing (CRISP). Figure 8 and Algorithm 1 demonstrate the main process of our method, which comprises five modules:
Path Generation
Given a question $q$, during each iteration, we generate new reasoning paths based on the existing prefix set $\mathcal{P}$:
$\mathcal{R}=\mathcal{R}\cup\mathcal{M}(q,n,\mathcal{P})$ (3)
In the generation process, the policy model generates $n$ complete sequences of remaining reasoning steps conditioned on $\mathcal{P}$ ($\mathcal{P}=\emptyset$ in the init iteration), rather than generating intermediate nodes step by step as in approaches like MCTS. This helps control the diversity of the search space and reduces the negative impact of excessive diversity on the reward model, as discussed in Cl.3.
State Aggregation
To further reduce the complexity of the state space and mitigate the impact of low-frequency negative examples on the reward model’s performance (as discussed in Cl.2), we define a final-answer-based state aggregation function $\psi$:
$\psi:\mathcal{R}\xrightarrow{}\mathcal{C}$ (4)
where $\mathcal{C}$ is the set of final answer clusters (i.e., all responses leading to the same answer), and for any path $r_{1},r_{2}\in\mathcal{R}$, we have:
$\psi(r_{1})=\psi(r_{2})\iff Answer(r_{1})=Answer(r_{2})$ (5)
All paths that produce the same final answer are mapped to the same cluster $\mathcal{C}_{j}\in\mathcal{C}$. As an example, in Module 2 of Figure 8, paths 1 and 3, both with the answer of -50, are assigned to the same cluster.
Reward Evaluation
After clustering the responses, we can convert the reward scores $f$ for each path into scores $\mathcal{F}$ for the corresponding clusters $\mathcal{C}_{j}$ (i.e., lines 17-20 in Algorithm 1):
$\mathcal{F}(\mathcal{C}_{j})=\sum_{x\in\mathcal{C}_{j}}f(x)$ (6)
In the implementation, we normalize $f(x)$ before summing. By additionally considering the frequency of the answers associated with each path during scoring, we can prevent the reward model from assigning excessively high scores to low-frequency responses, thereby mitigating the issue identified in Cl.2. We will later demonstrate the effectiveness of this clustering strategy through both ablation experiments (see §4.5) and theoretical analysis (see Appendix K).
Early Termination
This module controls when to exit the loop and return the final response. In addition to the standard exit condition of reaching the maximum number of iterations, we also control early termination by monitoring the number of clusters. If the number falls below a certain threshold (set to 2 in our work), it indicates that the question is relatively simple (as evidenced and discussed in Appendix E). In this case, the algorithm terminates, returning the answer corresponding to the most populated cluster, which is equivalent to SC. This not only reduces inference costs but also mitigates the issue of the reward model underperforming on simple questions (see Cl.1).
Prefix Extraction
In this module, we extract the top multiple prefixes as the new prefix set $\mathcal{P}$ for the next iteration, based on the scores of the paths and clusters. As illustrated in Module 5 of Figure 8, we first select the top-$k$ clusters with the highest scores (here, kk=1, so we select Cluster 1). Then, from the selected cluster(s), we choose the path with the highest score (in this case, 0.8 ¿ 0.7, so we select Path 3) to extract the prefix. Specifically, at the $i$-th generation, we extract the first $i$ steps of all paths as $\mathcal{P}$, and repeat the process until termination.
@@
Table 1: Accuracy comparison in main experiments, the best results are highlighted in bold.
Methods Qwen2.5-3B Llama3.1-8B
GSM8K MATH Olympiad GSM8K MATH Olympiad
CoT 0.78 0.46 0.24 0.85 0.38 0.11
Self-Consistency 0.83 0.64 0.31 0.91 0.57 0.16
Best-of-N + ORM 0.83 0.65 0.31 0.91 0.47 0.18
+ PRM 0.87 0.61 0.34 0.95 0.62 0.23
BoN Weighted + ORM 0.83 0.67 0.31 0.89 0.53 0.20
+ PRM 0.86 0.60 0.36 0.94 0.62 0.24
MCTS + ORM 0.92 0.67 0.34 0.90 0.43 0.13
+ PRM 0.95 0.71 0.31 0.95 0.57 0.19
Beam Search 0.95 0.73 0.34 0.94 0.56 0.15
Ours + ORM 0.91 0.70 0.36 0.89 0.49 0.18
+ PRM 0.96 0.76 0.39 0.95 0.67 0.26
@@ 4.2 Main Experiments
We compare the reasoning performance of our method with other advanced baselines, including: CoT (Wei et al., 2022), Self-Consistency (Wang et al., 2023), Best-of-N, BoN Weighted (Snell et al., 2024), MCTS (Hao et al., 2023) and Beam Search (Snell et al., 2024). For datasets, in addition to MATH-500 (Hendrycks et al., 2021; Lightman et al., 2024), we also validate our methods on GSM8K (Cobbe et al., 2021) and OlympiadBench (He et al., 2024). For models, we continue to select Qwen2.5-3B and Llama3.1-8B as the policy model, while using Skywork-Llama-3.1-8B (ORM) and Skywork-o1-PRM-Qwen-2.5-7B (PRM) as the reward model. We present more details in Appendix L.
@@
We demonstrate the result in Table 1, from which we can get the following conclusions: (1) Our proposed CRISP method significantly improves RM’s performance in reasoning tasks. Across all benchmarks and both model backbones, CRISP consistently outperforms existing RM-based inference approaches. Notably, on the Llama3.1-8B model, CRISP achieves a performance gain of up to 5.0% on the MATH dataset over the best-competing method. (2) The findings from the preceding analysis are reasonable. CRISP is specifically crafted to overcome the key issues of reward modeling revealed in §3. Its consistent and significant performance improvements provide strong empirical evidence that CRISP effectively mitigates these limitations, which are critical bottlenecks affecting the model’s reasoning performance. We present detailed experiments and discussions in Appendix O to further validate the stability and significance of the improvements achieved by our CRISP method.
4.3 Training-Time vs. Inference-Time Optimization
To demonstrate the continued necessity of our inference-time optimization approach amid the rising dominance of RL and SFT techniques represented by the DeepSeek-R1 series, we compare our method against the R1 model across different reasoning tasks, including math reasoning (MATH-500), commonsense reasoning (CSQA (Talmor et al., 2019)), social reasoning (SIQA (Sap et al., 2019)) and logical reasoning (LogiQA (Liu et al., 2020)). Specifically, given the same base model, we evaluate the accuracy and token consumption among its chat version (using CoT), the R1 distilled version, and our proposed method. From the results in Table 2, we can observe that: (1) Our method enables more efficient reasoning across all tasks. It achieves comparable reasoning tokens to the CoT method, while reducing output length by over 90% compared to the R1 model in the best case. (2) Our method exhibits stronger generalization capabilities. Although it underperforms the R1 model on math tasks, it consistently outperforms R1 on other reasoning benchmarks, with average gains of 10% and 5% accuracy across two backbones. This highlights the advantage of our inference-time optimization in generalizing across diverse scenarios.
Table 2: Comparison between R1 models and our method, the best accuracy are highlighted in bold.
Base Models Methods Math Commonsense Social Logical
Acc Length Acc Length Acc Length Acc Length
Qwen2.5-Math-1.5B Chat 0.52 1470 0.40 1400 0.46 1204 0.40 2790
R1-Distill 0.79 13421 0.47 6066 0.52 6407 0.35 12352
Ours 0.59 943 0.58 1004 0.61 1144 0.44 1143
Qwen2.5-Math-7B Chat 0.74 1855 0.58 1479 0.58 1388 0.49 2133
R1-Distill 0.88 9626 0.65 3612 0.66 2920 0.50 6492
Ours 0.79 987 0.72 1100 0.66 1059 0.59 2058
Figure 9: Performance comparison on other reasoning tasks (Llama3.1-8B + Skyworko1).
Figure 10: Performance comparison on other reward models (Llama3.1-8B on MATH).
4.4 Generalization Capability Evaluation
Results on More Tasks.
To ensure our method applies to tasks beyond mathematical reasoning, we introduce two additional tasks: logical reasoning (LogiQA (Liu et al., 2020)) and commonsense reasoning (CSQA (Talmor et al., 2019)), and compare the accuracy with other baselines on them. As shown in Figure 10, when using Llama3.1-8B as the policy model and Skyworko1 as the reward model, our method consistently outperforms all baselines across tasks, highlighting its versatility.
Results on More Reward Models.
To demonstrate the robustness of our method across different RMs, we further evaluate it using two additional RMs: Shepherd-Mistral-7B-PRM (Wang et al., 2024b) and Qwen2.5-Math-PRM-7B (Zhang et al., 2025b). We replicate the main experiment on the MATH dataset (200 samples) and report the result in Figure 10. The results show that our method still significantly outperforms other baselines when using other reward models. Even with a relatively weak reward model like Shepherd (achieving only 0.47 BoN performance), our method is able to maintain a high level of accuracy.
Table 3: Time cost comparison (s).
Dataset BoN MCTS Beam Ours
GSM8K 33.6 89.7 99.7 53.3
MATH 58.6 211.3 268.7 91.0
Table 4: Token consumption comparison.
Dataset BoN MCTS Beam Ours
GSM8K 6,340 9,282 8,828 3,499
MATH 11,550 18,014 27,012 11,535
4.5 Other Discussions
Cost Analysis
As an inference-time method, in addition to accuracy, reasoning cost is also an important factor to consider. We evaluate computational cost (token consumption and inference time) under consistent rollout numbers and device settings, with results demonstrated in Table 4 and Table 4. Our approach outperforms advanced RM-integrated methods such as MCTS and Beam Search in both time and token consumption across two datasets. Despite having a slightly higher inference time than BoN, our method offers an effective balance between efficiency and overall performance. We report the time-accuracy Compute-Return curve in Figure 30, which further substantiates this conclusion.
Ablation Study
We perform ablation experiments to validate the contribution of each module in the CRISP framework, with results summarized in Appendix M. The results show that removing any single module leads to a decline in performance. As our design is informed by the analysis presented in §3 (i.e., Cl.1-Cl.3), the results provide further empirical support for our findings.
5 Related Work
Inference-time Optimization Technique in LLM’s Reasoning
@@ 5 Related Work
Recent studies have demonstrated that large language models (LLMs) can be effectively enhanced through search-based optimization at inference time (OpenAI, 2024; Zeng et al., 2024; Zhao et al., 2024). These works primarily follow two approaches: optimizing the strategy for LLMs to search for answers (Hao et al., 2023; Snell et al., 2024; Bi et al., 2024; Qi et al., 2024) or improving the reward model’s ability to evaluate response quality (Wang et al., 2024b; Zhang et al., 2024; Setlur et al., 2024). However, most studies explore these two approaches separately, with limited research analyzing the impact of search factors on RM performance. Our work addresses this gap and proposes a new search strategy to mitigate RM’s deficiencies.
@@
The reward model plays a crucial role in complex reasoning tasks of LLMs (Zeng et al., 2024; Setlur et al., 2024; Wang et al., 2024b). Existing works mainly investigate the RM from two perspectives: evaluation and optimization. For the former, researchers design various datasets to evaluate the RM’s ability to distinguish between positive and negative responses (Lambert et al., 2024; Liu et al., 2024c; Zheng et al., 2024). For the latter, researchers focus on the training phase, improving the RM’s ability by synthesizing high-quality data (Wang et al., 2024b; Liu et al., 2024a) or optimizing the training algorithm (Zhang et al., 2024; Ankner et al., 2024; Lou et al., 2024). There is a lack of in-depth analysis of the potential issues RM faces during inference, as well as methods to optimize RM’s performance in the inference stage. Our work addresses the gaps left by these related studies.